OpenStudy (anonymous):

If the parent function f(x) = (2x − 3)3 is transformed to g(x) = (-2x + 3)3, which type of transformation occurs? PLEASE HELP! :/

OpenStudy (anonymous):

\(-(2x-3)^3=(-1)^3(2x-3)^3=(-1\cdot(2x-3))^3=(-2x+3)^3\) so \(g(x)=-f(x)\)

OpenStudy (anonymous):

this is a reflection about the x-axis as \((x,y)\mapsto(x,-y)\)